OpenStudy (anonymous):

22. Solve. 8/x + 15/x^2 = -1

OpenStudy (anonymous):

Okay, first simplify the equation: \[\frac{ 8x + 15 }{ x ^{2}} = -1\] \[8x + 15 = -x ^{2}\] \[x ^{2} + 8x + 15 = 0\] Can you solve it now?

OpenStudy (anonymous):

No.. im sorry but im still lost

OpenStudy (anonymous):

That's okay, you can either solve it through the quadratic formula method or the splitting method: The equation can be written as : \[x ^{2} +5x + 3x + 15 = 0\] \[x(x+5) + 3(x+5)\] \[(x+5) (x+3)\] Therefore, x = -5 or x = - 3 Get it now? :-}

OpenStudy (anonymous):

Yess I understand now. Thanks a lot (: